Adv. Math. Econ. 9, 117-126 (2006)

The distribution of firm size

Na Zhang

Graduate School of Mathematical Sciences, The University of Tokyo, 3-8-1 Komaba, Meguro-ku, Tokyo, 153-8914, Japan Received: January 20, 2005 Revised: September 5, 2005 JEL classification: C0 Mathematical Subject Classification (2000): 60E07, 60F99 Abstract. In this paper, we study the distribution of firm size by using a model based on Sato’s paper in 1970, and proved the static distribution of firm size satisfies Pareto distribution in its upper tail. Key words: firm size, Pareto distribution, Yule distribution